OpenStudy (anonymous):

why does the limit of (x+1)^2 -2 / x as x approaches 0 not exist? (x cannot equal 0)

OpenStudy (anonymous):

why does \[\lim \frac{ (x+1)^2 -2}{ x }\] as x ->0 not exist?

OpenStudy (anonymous):

(x^2+2x+1-2)/x (x^2+2x-1)/x x+2-1/x x => 0+ => 0+2-inff =>-inf x => 0- => 0+2+inf =>+inf so limit does not exist

OpenStudy (anonymous):

how did u get x+2-1 (2nd to 3rd step)?

OpenStudy (agent0smith):

Plug in x=0 and you'll see why (you'll get -1/0) There's no need for algebraic manipulation... if a limit approaches pos or neg infinity, then the limit does not exist.

OpenStudy (agent0smith):

@sara17 it doesn't matter that the left and right hand limits aren't equal in this case. Neither the left or right limits exist since they both approach infinity.

OpenStudy (anonymous):

the question is : Let f(x) de defined by f(x)=\[\lim \frac{ (x+1)^2 -2 }{ x }\] for \[x \neq 0\]. Explain why the limit of this function as x approaches o does not exist

OpenStudy (agent0smith):

Well then, that's a diff question...

OpenStudy (anonymous):

yup, when u answered the question, I remember to include that detail. Really sorry :S

OpenStudy (agent0smith):

btw, "how did u get x+2-1 (2nd to 3rd step)?" from this: (x^2+2x-1)/x x+2-1/x comes from this: \[\Large \frac{ x^2 +2x -1 }{ x } = \frac{ x^2 }{ x } + \frac{2x }{ x } - \frac{1 }{ x }\]

OpenStudy (zzr0ck3r):

if you approach from the left you get positive infinity, and vice versa

OpenStudy (zzr0ck3r):

are you talking about extended real numbers?

OpenStudy (anonymous):

no, but I understand what agent0smith and sara17 were saying now. Thank you all!

OpenStudy (anonymous):

actually, maybe it is all real numbers. Not too sure